Consider two thin, concentric conducting spherical shells with radii r1 = 0.50 m and r2 = 1.0 m. A charge of +1.0 mC is placed on the inner sphere and a charge of +2.0 mC is placed on the outer sphere. Set the potential at infinity to be 0. Determine (a) the field inside the inner sphere, (b) the charge on the inner surface of the outer conductor, (c) the magnitude of the E-field midway between the inner and outer spheres, (d) the field at the outer surface of the outer sphere, and (e) the electrical potential at a distance r = 2.0 m from the center of the two spheres.
Answer #1

The electric field and electric potential

using Gausss law thome rir qen =0 LE=0 mi 9in= -9 = -1.0mc & Eda = gem E(4102) = que 6 = 91 4TTEor 2 - 9X100 x 1x103 TOI +12Е (41 к2 ) = 9, +1, fo С — 9, +42 , Т- с эхiox зxis 3 о2. E = 2.781 07 Nich e E = 9, +92 af v- - Sear 4 г. x- Ү-- 4, +4. — 9